OpenStudy (jakyfraze):

someone please help me The spoke of a wheel reaches from the center of the wheel to its rim. If the circumference of the rim of the wheel is 27 inches, how long is each spoke? Round your answer to the nearest hundredth. A. 2.07 B. 2.93 C. 4.30 D. 8.59

OpenStudy (jakyfraze):

@Jazzy17999 can you please help me?

OpenStudy (anonymous):

ok hold on a sec

OpenStudy (jakyfraze):

@Jazzy17999 thank you

OpenStudy (anonymous):

ur welcome

OpenStudy (jakyfraze):

@carlyleukhardt can you help me with this equation? i got the answer D. then o got the answer C. please help me

OpenStudy (carlyleukhardt):

The answer is C

OpenStudy (anonymous):

Yes the answer is C

OpenStudy (anonymous):

I was trying to give you the reason why its c

OpenStudy (jakyfraze):

please do @Jazzy17999

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

C=2PI*r r=C/2PI=27/6.28 inches 2pi = 6.28 so, you have 2pi r = C 6.28r = 27 now just divide by the 6.28 to get r. When you divide 27 by 6.28 you get 4.29936305732 rounded is 4.30 As I said before R is C (4.30)

OpenStudy (anonymous):

Hope this helps u understand

OpenStudy (jakyfraze):

thank you so much!!! but how is R=C?

OpenStudy (anonymous):

C is 4.30 so R is 4.30 or C

OpenStudy (jakyfraze):

i don't get in. sorry

OpenStudy (anonymous):

ok The letter for the answer 4.30 is C

OpenStudy (anonymous):

C. 4.30

OpenStudy (jakyfraze):

oh okay lol sorry

OpenStudy (jakyfraze):

@Jazzy17999 can you help me with this please? Which one of the following statements is true of perpendicular lines? A. They intersect to form an obtuse angle. B. They intersect at one point. C. They intersect to form an acute angle. D. They never intersect.

OpenStudy (anonymous):

b

OpenStudy (anonymous):

@jakyfraze

OpenStudy (jakyfraze):

oh my god thank you so much @Jazzy17999 :D

OpenStudy (anonymous):

lol ur welcome and also thnx for the medal

OpenStudy (jakyfraze):

@ your welcome :D

OpenStudy (jakyfraze):

@Jazzy17999 one more question please

OpenStudy (jakyfraze):

What is the circumference of a circle with a diameter of 30 centimeters? A. 94.2 centimeters B. 188.4 centimeters C. 9.55 centimeters D. 47.1 centimeters

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Do u know how to find the circumference of a circle?

OpenStudy (anonymous):

@jakyfraze

OpenStudy (jakyfraze):

@Jazzy17999 im not sure. can you sure me please? 30 x 3.14?

OpenStudy (anonymous):

Yes that is how @jakyfraze you multiply 30 by 3.14 in this case to find the radius.

OpenStudy (jakyfraze):

@Jazzy17999 but some questions you have to use pi2= 6.28 or pi= 3.14

OpenStudy (anonymous):

yup

OpenStudy (jakyfraze):

@Jazzy17999 why is that? please tell me. I'm really confuse

OpenStudy (anonymous):

pi2 is 2x 3.14=6.28

OpenStudy (anonymous):

What grade r u in?

OpenStudy (anonymous):

7th or 8th?

OpenStudy (anonymous):

btw if u want to learn more about circumference http://www.mathgoodies.com/lessons/vol2/circumference.html here's a good website. @jakyfraze

OpenStudy (jakyfraze):

yeah i know that but some questions i did i didn't have to you pi2= 6.28, i had to use 3.14

OpenStudy (anonymous):

That was for a specific problem that u did it's not for every problem.

OpenStudy (anonymous):

i have to go but, i hope all that helped.

OpenStudy (jakyfraze):

@Jazzy17999 ohh okay that make sense lol. thank you for everything :)

OpenStudy (anonymous):

Your welcome very much!!! @jakyfraze